OpenStudy (anonymous):

use the limit process to prove the derivative of the sine function. use the quotient rule to prove the derivative of the tangent function.

OpenStudy (larseighner):

For this you need: \[\large \lim_{\theta\rightarrow0} {\frac{\cos\theta - 1}{\theta}} = 0\]

OpenStudy (anonymous):

I use that for the sine function?

OpenStudy (larseighner):

Set up your difference quotient and see if you don't.

OpenStudy (larseighner):

You do also need \[ \large \lim_{\theta\rightarrow0} {\sin\theta \over \theta} = 1 \] Since one of these can be derived of the other, the proof of them is generally offered before trig derivatives.